OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? @Mathematics

OpenStudy (anonymous):

Tim is 9 years old.

OpenStudy (anonymous):

that can't be right?!

OpenStudy (anonymous):

Sorry. I got something wrong in my working. :P

OpenStudy (anonymous):

tim is 19. i think :P

OpenStudy (anonymous):

nonononono, he is 13. sorry :P

OpenStudy (anonymous):

Nonono. He's 13!

OpenStudy (anonymous):

t+10=j j-3=2(t-3) t+7=2t-6 t=13

OpenStudy (anonymous):

Tim is 13.

OpenStudy (anonymous):

Tim + 10 years _ coach johnson. j-3= 2(t-3) t+7=2t-6 subtract 2t from both sides. you end up with t= 6+7 so, t=13